# Maximum Inner Product is Query-Scaled Nearest Neighbor

### v1

@@ 8. Conclusion
In this paper, we align Maximum Inner Product Search (MIPS) with Nearest Neighbor Search (NNS) through robust theoretical foundations, presenting the first analytical framework for $\mathsf{MIPS}$ efficiency on graph indices. To address the inherent challenges of biased solutions and excessive explorations in graph-based $\mathsf{MIPS}$, we propose a practical framework, featuring a novel graph index, $\mathsf{PSP}$, and a novel lightweight adaptive early termination mechanism, AET. Extensive experiments validate the theoretical soundness, efficiency, scalability, and practical applicability of our approaches. Notably, our method has been deployed in the search engine of the Shopee online shopping platform, demonstrating superior performance. Additionally, we contribute the Commerce100M dataset to the public community, addressing the lack of e-commerce data in this domain.

### v2

@@ 8. Conclusion
In this paper, we align Maximum Inner Product Search (MIPS) with Nearest Neighbor Search (NNS) through robust theoretical foundations, presenting the first analytical framework for $\mathsf{MIPS}$ efficiency on graph indices. To address the inherent challenges of biased solutions and excessive explorations in graph-based $\mathsf{MIPS}$, we propose a practical framework, featuring a novel graph index, $\mathsf{PSP}$, and a novel lightweight adaptive early termination mechanism, AET. Extensive experiments validate the theoretical soundness, efficiency, scalability, and practical applicability of our approaches. Notably, our method has been deployed in the Shopee search engine, demonstrating superior performance. Additionally, we contribute the Commerce100M dataset to the public community, addressing the lack of e-commerce data in this domain.
Acknowledgements.
This work was supported in part by the NSFC under Grants No. (62025206 and U23A20296), Zhejiang Province’s “Lingyan” R&D Project under Grant No. 2024C01259, CCF-Aliyun2024004, Ningbo Yongjiang Talent Introduction Programme (2022A-237-G). Cong Fu is the main project advisor. Xiangyu Ke is the corresponding author of this work.